# UltraClean: A Simple Framework to Train Robust Neural Networks against Backdoor Attacks

### v1

@@ C.2 Clean-Label Attacks
C.2.1 SIG
The ablation study of baseline images generated by different denoising functions and $\beta$ for detection on the poisoned class and the whole dataset are summarized in Tables 11 and 12, respectively. For the detection on the poisoned class, using the single “mean” baseline images achieves a slightly higher detection rate than using the single “median” function and the aggregation of mean and median for most classes. Results of the detection on the whole dataset reveal the same trend. We then compare the post-clean ASR and present the results in Figure 3. Although using the single “mean” baselines perform better in detection rate, there is still a considerable performance gap to applying both baseline images in most cases. In general, the aggregation of median and mean denoising shows the best performance in degrading the post-clean ASR. We believe this is because the aggregation leads to a larger noise difference, which makes poisoned samples easier to detect. The ablation study of $\beta$ shows that UltraClean’s performance against SIG is sensitive to the adjustment of the removal threshold. A larger $\beta$ can detect more poisoned samples and better reduce ASR.
@@ C.2.1 SIG
C.2.2 LCBD
Tables 13 and 14 present the detection performance against LCBD on the poisoned class and the whole dataset, respectively. It can be seen that the single “median” denoised baseline images outperform the single “mean” denoised baseline images by a large margin in detecting poisoned samples formed by LCBD, which is different from the SIG attacks. On the other hand, the aggregation of median and mean denoising still has a comparable detection rate to the single “median” denoising. This trend is even prominent for the detection on the whole dataset where employing the single “mean” denoising images can barely detect any poisoned sample. Meanwhile, we found the post-clean ASR of employing the single “median” and the aggregation version are close, since they detect almost the same poisoned samples. In contrast to SIG, defending against LCBD is not too sensitive to the selection of $\beta$. As can be seen in Table 14, UltraClean still detects 88.80% poisoned samples when $\beta$ is small.
@@ C.2.2 LCBD
C.2.3 HTBD
Unlike SIG and LCBD, for the detection on the poisoned class, HTBD attacks are sensitive to the single “median” and the single “mean” denoised baseline images. However, the aggregation version consistently achieves the highest detection rate for all target classes, as shown in Table 15. We can observe from Figure 4 that the performance of the post-clean ASR is similar to the detection rate, where the aggregation version always performs the best. For the detection on the whole dataset, applying both the aggregation version and the single “median” denoised baseline images detect considerable poisoned samples while applying the single “mean” denoising only detects a few of them. UltraClean is resistant to the selection of $\beta$ in defending against HTBD, only a removal threshold of 5% is capable of entirely mitigating HTBD from models.
@@ C.2.3 HTBD
Appendix D Robustness against Adaptive Attacks
D.1 Adaptive Attacks of Dirty-Label Attacks
We first evaluate UltraClean’s performance against adaptive attacks of dirty-label attacks. In order to bypass the defense of UltraClean, attackers can reduce backdoor trigger size or trigger blended/transparency ratio to craft harder-to-detect poisoned samples. Note that attackers tend not to reduce the poisoning ratio for dirty-label attacks since it is usually below 10% to evade label sanitization defense [53]. A lower poisoning ratio may render a failure of attacks. We conduct the adaptive attacks by adjusting the trigger size ($\mathcal{T}_{s}$) of BadNets from $1\times 1$ to $3\times 3$, the trigger transparency ($\mathcal{T}_{t}$) of Trojan from 0.1 to 0.5 and blended ratio ($\alpha$) of Blended from 0.1 to 0.3. The results are summarized in Table 10. We can see that while most attacks can achieve high ASR even with harder-to-detect triggers, UltraClean demonstrates exceptional robustness to various adaptive attacks. UltraClean identifies at least 92% poisoned samples for almost all adaptive attacks and successfully mitigates backdoors in all models. The only exception is the Trojan attack with $\mathcal{T}_{t}=0.1$, where the attack achieves 9.43% ASR even without a defense. Although UltraClean only detects 32% of the poisoned data, it still completely remove the backdoor. The experiment results validates UltraClean’s robustness against dirty-label adaptive attacks.
@@ D.1 Adaptive Attacks of Dirty-Label Attacks
D.2 Adaptive Attacks of Clean-Label Attacks
D.2.1 LCBD
In the main manuscript, we extensively evaluated UltraClean against LCBD with GAN-based and AE-based approaches. UltraClean has shown superior performance in detecting and mitigating the LCBD backdoor. Here, we further study the effectiveness of UltraClean against adaptive attacks. Attackers may change poisoning ratio and trigger amplitude to make LCBD more stealthy. In the previous experiments, we injected 40% poisoned sample into the target class. Although the poisoning ratio (i.e., the fraction of poisoned data injected into the training dataset) over the entire dataset is small, 40% is a relatively high poisoning ratio over the poisoned class. The original LCBD paper [62] shows that with lower poisoning ratios (lower poisoning ratios usually render higher stealthiness), it can still achieve a high attack success rate. Thus, we investigate if UltraClean is still effective with lower poisoning ratios. We perform the experiment with the setting “AE ($\ell_{2}$, $\epsilon$ = 1200)” since it achieves the best attack result. As shown in Table 18, even with 10% poisoned data in the target class (only 1% poisoning ratio over the entire training dataset), LCBD still achieves $\sim$99ASR. Even under this circumstance, UltraClean still demonstrates a stable detection performance, reaching at least a 95% detection rate and reducing the ASR by over 98%. Another critical factor to make the attack more stealthy is the trigger amplitude. Smaller amplitude means less trigger visibility. In the previous experiments, we set the amplitude to 64 to secure the attack success rate. Here we reduce the amplitude to 16 and 32 to examine if UltraClean can still detect poisoned samples. The experiment results are summarized in the last two rows in Table 18. We find that UltraClean still detects at least 90% poisoned samples, even with an extremely low trigger amplitude. Note that with a 16 trigger amplitude, the attack has failed to inject an effective backdoor since the pre-clean ASR is only 1.12%. The results further demonstrate the effectiveness of UltraClean against different attack settings, indicating its robustness in protecting neural networks against various adaptive attacks.
@@ D.2.1 LCBD
$\epsilon$=16 $\mathcal{T}_{s}$=30 Class ID
@@
$\epsilon$=8 $\mathcal{T}_{s}$=30 Class ID
@@
$\epsilon$=32 $\mathcal{T}_{s}$=30 Class ID
@@
$\epsilon$=16 $\mathcal{T}_{s}$=15 Class ID
@@
$\epsilon$=16 $\mathcal{T}_{s}$=60 Class ID
@@
D.2.2 HTBD
For HTBD attacks, there are two crucial parameters: the maximum perturbation $\epsilon$ and the trigger size ($\mathcal{T}_{s}$), which affect the performance and stealthiness of poisoned samples. Attackers may try to evade or break (i.e., achieve a higher ASR in the presence of detection) the defense of UltraClean by adjusting these parameters. Here we vary these two parameters to study their effects on UltraClean. Results are presented in Tables 19 and  20. Recall that HTBD generates poisoned samples by adding perturbations to poisoned images to match source images’ representations in the feature space. Therefore, $\epsilon$ in the HTBD is akin to the adversarial perturbations of the LCBD, which is the upper bound of perturbation imposed on the image. The trigger patched on source images affects their representations in the feature space. Thus the trigger size is also an important parameter for poisoned data generation. According to [57], a larger trigger size means larger perturbations added on poisoned images and always leads to better attack efficiency. As shown in the results, UltraClean shows highly consistent performance against all adaptive attacks of different perturbation and trigger size. In all experiments, UltraClean detects almost all poisoned samples and reduces the post-clean ASR to 0% in most cases.
@@ Appendix E Comparison to Other SOTA Defenses
We have shown that UltraClean outperforms STRIP and SVD, which are poison-filtering-based defenses. To better demonstrate the effectiveness of UltraClean, we compare it to other SOTA defenses ABL [22] and ANP [20], which are not poison-filtering-based. Results are summarized in Table 21. ABL is a poison suppression-based defense that eliminates backdoors by unlearning the poisoned samples in the dataset. ANP is a model reconstruction-based defense that prunes some sensitive neurons to purify the injected backdoor. We conduct the experiment against BadNets on CIFAR-10. It can be seen that UltraClean achieves comparable clean accuracy and ASR while cleansing the dataset for reusability, as opposed to ABL and ANP that do not remove poisoned data.
@@ Appendix F Effectiveness against Sleeper Agent
We evaluate the effectiveness of UltraClean against various prestigious attacks. We notice the recently proposed clean-label backdoor attack Sleeper Agent (SA) [21] employs a different attack mechanism from previous attacks, which crafts poisoned samples via gradient matching [25]. We evaluate UltraClean’s performance against this latest type of attack in Table 22. We carry out the experiment on CIFAR-10 using ResNet-18. It can be seen that UltraClean still significantly reduces the ASR and successfully mitigate the backdoor effect, which again validates that UltraClean is effective against both clean and dirty-label attacks regardless of attack mechanisms.
@@
Sleeper Agent [21] 92.31% 85.27% -

### v2

@@ C.2 Clean-Label Attacks
SIG. The ablation study of baseline images generated by different denoising functions and $\beta$ for detection on the poisoned class and the whole dataset are summarized in Tables 11 and 12, respectively. For the detection on the poisoned class, using the single “mean” baseline images achieves a slightly higher detection rate than using the single “median” function and the aggregation of mean and median for most classes. Results of the detection on the whole dataset reveal the same trend. We then compare the post-clean ASR and present the results in Figure 4. Although using the single “mean” baselines perform better in detection rate, there is still a considerable performance gap to applying both baseline images in most cases. In general, the aggregation of median and mean denoising shows the best performance in degrading the post-clean ASR. We believe this is because the aggregation leads to a larger noise difference, which makes poisoned samples easier to detect. The ablation study of $\beta$ shows that UltraClean’s performance against SIG is sensitive to the adjustment of the removal threshold. A larger $\beta$ can detect more poisoned samples and better reduce ASR.
@@
LCBD. Tables 13 and 14 present the detection performance against LCBD on the poisoned class and the whole dataset, respectively. It can be seen that the single “median” denoised baseline images outperform the single “mean” denoised baseline images by a large margin in detecting poisoned samples formed by LCBD, which is different from the SIG attacks. On the other hand, the aggregation of median and mean denoising still has a comparable detection rate to the single “median” denoising. This trend is even prominent for the detection on the whole dataset where employing the single “mean” denoising images can barely detect any poisoned sample. Meanwhile, we found the post-clean ASR of employing the single “median” and the aggregation version are close, since they detect almost the same poisoned samples. In contrast to SIG, defending against LCBD is not too sensitive to the selection of $\beta$. As can be seen in Table 14, UltraClean still detects 88.80% poisoned samples when $\beta$ is small.
@@
HTBD. Unlike SIG and LCBD, for the detection on the poisoned class, HTBD attacks are sensitive to the single “median” and the single “mean” denoised baseline images. However, the aggregation version consistently achieves the highest detection rate for all target classes, as shown in Table 15. We can observe from Figure 5 that the performance of the post-clean ASR is similar to the detection rate, where the aggregation version always performs the best. For the detection on the whole dataset, applying both the aggregation version and the single “median” denoised baseline images detect considerable poisoned samples while applying the single “mean” denoising only detects a few of them. UltraClean is resistant to the selection of $\beta$ in defending against HTBD, only a removal threshold of 5% is capable of entirely mitigating HTBD from models.
@@
Appendix D Robustness against Adaptive Attacks
D.1 Adaptive Attacks of Dirty-Label Attacks
We first evaluate UltraClean’s performance against adaptive attacks of dirty-label attacks. In order to bypass the defense of UltraClean, attackers can reduce backdoor trigger size or trigger blended/transparency ratio to craft harder-to-detect poisoned samples. Note that attackers tend not to reduce the poisoning ratio for dirty-label attacks since it is usually below 10% to evade label sanitization defense [DBLP:journals/corr/abs-1803-00992]. A lower poisoning ratio may render a failure of attacks. We conduct the adaptive attacks by adjusting the trigger size ($\mathcal{T}_{s}$) of BadNets from $1\times 1$ to $3\times 3$, the trigger transparency ($\mathcal{T}_{t}$) of Trojan from 0.1 to 0.5 and blended ratio ($\alpha$) of Blended from 0.1 to 0.3. The results are summarized in Table 10. We can see that while most attacks can achieve high ASR even with harder-to-detect triggers, UltraClean demonstrates exceptional robustness to various adaptive attacks. UltraClean identifies at least 92% poisoned samples for almost all adaptive attacks and successfully mitigates backdoors in all models. The only exception is the Trojan attack with $\mathcal{T}_{t}=0.1$, where the attack achieves 9.43% ASR even without a defense. Although UltraClean only detects 32% of the poisoned data, it still completely remove the backdoor. The experiment results validates UltraClean’s robustness against dirty-label adaptive attacks.
@@ D.1 Adaptive Attacks of Dirty-Label Attacks
D.2 Adaptive Attacks of Clean-Label Attacks
LCBD. In the main manuscript, we extensively evaluated UltraClean against LCBD with GAN-based and AE-based approaches. UltraClean has shown superior performance in detecting and mitigating the LCBD backdoor. Here, we further study the effectiveness of UltraClean against adaptive attacks. Attackers may change poisoning ratio and trigger amplitude to make LCBD more stealthy. In the previous experiments, we injected 40% poisoned sample into the target class. Although the poisoning ratio (i.e., the fraction of poisoned data injected into the training dataset) over the entire dataset is small, 40% is a relatively high poisoning ratio over the poisoned class. The original LCBD paper [DBLP:journals/corr/abs-1912-02771] shows that with lower poisoning ratios (lower poisoning ratios usually render higher stealthiness), it can still achieve a high attack success rate. Thus, we investigate if UltraClean is still effective with lower poisoning ratios. We perform the experiment with the setting “AE ($\ell_{2}$, $\epsilon$ = 1200)” since it achieves the best attack result. As shown in Table 18, even with 10% poisoned data in the target class (only 1% poisoning ratio over the entire training dataset), LCBD still achieves $\sim$99% ASR. Even under this circumstance, UltraClean still demonstrates a stable detection performance, reaching at least a 95% detection rate and reducing the ASR by over 98%. Another critical factor to make the attack more stealthy is the trigger amplitude. Smaller amplitude means less trigger visibility. In the previous experiments, we set the amplitude to 64 to secure the attack success rate. Here we reduce the amplitude to 16 and 32 to examine if UltraClean can still detect poisoned samples. The experiment results are summarized in the last two rows in Table 18. We find that UltraClean still detects at least 90% poisoned samples, even with an extremely low trigger amplitude. Note that with a 16 trigger amplitude, the attack has failed to inject an effective backdoor since the pre-clean ASR is only 1.12%. The results further demonstrate the effectiveness of UltraClean against different attack settings, indicating its robustness in protecting neural networks against various adaptive attacks.
$\epsilon$=16
$\mathcal{T}_{s}$=30
Class ID
@@ D.2 Adaptive Attacks of Clean-Label Attacks
$\epsilon$=8
$\mathcal{T}_{s}$=30
Class ID
@@
$\epsilon$=32
$\mathcal{T}_{s}$=30
Class ID
@@
$\epsilon$=16
$\mathcal{T}_{s}$=15
Class ID
@@
$\epsilon$=16
$\mathcal{T}_{s}$=60
Class ID
@@
HTBD. For HTBD attacks, there are two crucial parameters: the maximum perturbation $\epsilon$ and the trigger size ($\mathcal{T}_{s}$), which affect the performance and stealthiness of poisoned samples. Attackers may try to evade or break (i.e., achieve a higher ASR in the presence of detection) the defense of UltraClean by adjusting these parameters. Here we vary these two parameters to study their effects on UltraClean. Results are presented in Tables 19 and  20. Recall that HTBD generates poisoned samples by adding perturbations to poisoned images to match source images’ representations in the feature space. Therefore, $\epsilon$ in the HTBD is akin to the adversarial perturbations of the LCBD, which is the upper bound of perturbation imposed on the image. The trigger patched on source images affects their representations in the feature space. Thus the trigger size is also an important parameter for poisoned data generation. According to [DBLP:conf/aaai/SahaSP20], a larger trigger size means larger perturbations added on poisoned images and always leads to better attack efficiency. As shown in the results, UltraClean shows highly consistent performance against all adaptive attacks of different perturbation and trigger size. In all experiments, UltraClean detects almost all poisoned samples and reduces the post-clean ASR to 0% in most cases.
@@ Appendix E Comparison to Other SOTA Defenses
We have shown that UltraClean outperforms STRIP and SVD, which are poison-filtering-based defenses. To better demonstrate the effectiveness of UltraClean, we compare it to other SOTA defenses ABL [DBLP:journals/corr/abs-2110-11571] and ANP [wu2021adversarial], which are not poison-filtering-based. Results are summarized in Table 21. ABL is a poison suppression-based defense that eliminates backdoors by unlearning the poisoned samples in the dataset. ANP is a model reconstruction-based defense that prunes some sensitive neurons to purify the injected backdoor. We conduct the experiment against BadNets on CIFAR-10. It can be seen that UltraClean achieves comparable clean accuracy and ASR while cleansing the dataset for reusability, as opposed to ABL and ANP that do not remove poisoned data.
@@ Appendix F Effectiveness against Sleeper Agent
We evaluate the effectiveness of UltraClean against various prestigious attacks. We notice the recently proposed clean-label backdoor attack Sleeper Agent (SA) [souri2022sleeper] employs a different attack mechanism from previous attacks, which crafts poisoned samples via gradient matching [DBLP:conf/iclr/GeipingFHCT0G21]. We evaluate UltraClean’s performance against this latest type of attack in Table 22. We carry out the experiment on CIFAR-10 using ResNet-18. It can be seen that UltraClean still significantly reduces the ASR and successfully mitigate the backdoor effect, which again validates that UltraClean is effective against both clean and dirty-label attacks regardless of attack mechanisms.
@@
Sleeper Agent [souri2022sleeper] 92.31% 85.27% -